OpenStudy (anonymous):

for which equatuation is x=4 a solution a. 5(3-2x)=7 b. -3(x-2)=-6 c. 2(x-5)=3 d. 4(3x+1)=49

OpenStudy (anonymous):

equation* whoopsie.

OpenStudy (radar):

b. might do it.

OpenStudy (radar):

-3(4-2)=-6 -3(2)=-6 -6=-6 By Jove thats is it.